Question 3 Solve for the open circuit voltage drop from a to b in Circuit 9 ΚΩ 4 ΚΩ 6 ΚΩ 3 2 mA 8 V. 8ΚΩ 13.6V 132 13 ν Π
Answer #1

Answer:13.2V

Redrawing the given network:

Clearly V2= 8

Now, by nodal analysis

\frac{V_{1}-V_{3}}{9}+\frac{V_{1}-8}{4}=2\Rightarrow 0.3611V_{1}-0.1111V_{3}=4

\frac{V_{3}-V_{1}}{9}+\frac{V_{3}-8}{6}+\frac{V_{3}}{8}=0\Rightarrow -0.1111V_{1}+0.4027V_{3}=1.3333

by solving the above two equations

V_{1}=13.21V

V_{3}=6.95V

Hence open-circuit voltage drop from a to b is 13.21